logIVU H₂O concentration of HCl is 2.0 x 100 M, What is the pH
Answer #1

briven that Hat H20 - 12ot tu- A [H3O+] = 2x10-10 ...pH = -log[Hgot] = -log (2x10) = 9.699 . Required /pH = 9.7